h(m) sand 13 1 04 3.3 use total Sand stresses to deter mine . rave rave Depth u
Answer #1

Solution:

a) Vertical Stress (\sigma _{v}):

For the active case,

at 1m depth: (dry-sand) 1 = 17 * 1 = 17 kN/m2

at 3m depth: (dry-sand) * = 17 * 3 = 51 kN/m2

at 6m depth: (dry-sand) *3sat-gravel = 17 * 3 +23 * 3 = 120 kN/m2

For the passive case,

at 3m depth: 0

at 6m depth: sat-gravel = 23 * 3 = 69 kN/m2

b) Pore pressure (u):

For the active case,

at 1m depth: W * = 0 kN/m2

at 3m depth: W * = 0 kN/m2

at 6m depth: W * = 9.81 * 3 = 29.43 kN/m2

For the passive case,

at 3m depth: 0

at 6m depth: W * = 9.81 * 3 = 29.43 kN/m2

c) Effective vertical Stress (\sigma _{v}'):

For the active case,

at 1m depth: \sigma _{v}-u   = 17 - 0 = 17 kN/m2

at 3m depth: \sigma _{v}-u = 51 - 0 = 51 kN/m2

at 6m depth: \sigma _{v}-u = 120 - 29.43 = 90.57 kN/m2

For the passive case,

at 3m depth: \sigma _{v}-u = 0

at 6m depth: \sigma _{v}-u   = 69 - 29.43 = 39.57 kN/m2

d) Effective horizontal Stress (\sigma _{H}'):

For the active case,

at 1m depth: \sigma _{v}*K_{a}   = 17 * 0.4 = 6.8 kN/m2

at 3m depth (just above the junction): \sigma _{v}*K_{a} = 51 * 0.4 = 20.40 kN/m2

at 3m depth (just below the junction): \sigma _{v}*K_{a}   = 51 * 0.3 = 15.30 kN/m2

at 6m depth: \sigma _{v}*K_{a} = 90.57 * 0.3 = 27.17 kN/m2

For the passive case,

at 3m depth: \sigma _{v}*K_{p} = 0

at 6m depth: \sigma _{v}*K_{p}   = 3.6 * 39.57 = 142.45 kN/m2

e) Horizontal Stress (\sigma _{H}):

For the active case,

at 1m depth: \sigma _{v}'*K_{a}   = 17 * 0.4 = 6.8 kN/m2

at 3m depth (just above the junction): \sigma _{v}'*K_{a} = 51 * 0.4 = 20.40 kN/m2

at 3m depth (just below the junction): \sigma _{v}'*K_{a} = 51 * 0.3 = 15.30 kN/m2

at 6m depth: \sigma _{v}'*K_{a}+\gamma _{w}*H = 90.57 * 0.3 + 9.81 * 3 = 57.14 kN/m2

For the passive case,

at 3m depth: \sigma _{v}'*K_{p} = 0

at 6m depth: \sigma _{v}'*K_{p}+\gamma _{w}*H   = 3.6 * 39.57 + 9.81 * 3 = 171.88 kN/m2